Consider the pdf of a Normal variable 𝑿 is defined as "f(x)= (1\/\u221a(98pi)^e^-((x+11)^2\/98)"

i) Find the mgf of 𝑿.

ii) Evaluate 𝑷(βˆ’πŸ‘ > βˆ’π‘Ώ > πŸπŸ‘).

iii) Find the value of π‘ͺ such that 𝑷(|𝑿 + 𝟏𝟏| β‰₯ π‘ͺ) = 𝟎. πŸŽπŸ”πŸπŸ’.

iv) Find βˆ’π’πŸŽ.πŸŽπŸŽπŸ‘πŸ and convert to 𝑿.

1
Expert's answer
2021-10-01T16:25:17-0400

i)


"N(x; \mu,\sigma^2) =\dfrac{1}{\sqrt{2\pi\sigma^2}} 
e^{-{1 \over 2}(x-\mu)^2/\sigma^2}"

Then "\mu=-11, \sigma^2=49."

The moment generating function corresponding to the normal probability density function"N(x; \u00b5, \u03c3^2)" is the function "M_{x}(t) = exp\{{\mu t + σ^2t ^2/2}\}"


"M_{x}(t) = exp\{{-11 t + 49t ^2/2}\}"

ii)


"P(-3>-X>13)=P(-13<X<3)"

"=P(X<3)-P(X\leq-13)"

"=P(Z<\dfrac{3-(-11)}{7})-P(Z\leq\dfrac{-13-(-11)}{7})"

"=P(Z<2)-P(Z\leq-0.2857)"

"=0.9772499-0.3875485=0.589701"

iii)

"P(|X+C|\geq11)=0.0614"

"P(X\leq-11-C)=0.0307"

"P(Z\leq\dfrac{-11-C-(-11)}{7})=0.0307"

"P(Z\leq\dfrac{-C}{7})=0.0307"

"\dfrac{-C}{7}=-1.870604"

"C=13.094228"


iv)


"P(Z\leq z)=0.0031"

"z=-2.737012"

"-z=2.737012"

"\dfrac{x-(-11)}{7}=-2.737012"

"x=-30.159"

"-x=30.159"